# OmniAD: View-Invariant Unsupervised Scene Pixelwise Anomaly Detection with Adaptive View Synthesis

### v2

@@ Abstract
Visual anomaly detection in the built environment can be a valuable tool for applications such as infrastructure assessment, construction monitoring, security surveillance, and urban planning. Anomaly detection approaches are typically unsupervised and work by detecting deviations from an expected state where no assumptions are made exact type of deviation. Unsupervised pixel-level anomaly detection methods have been developed to successfully recognize and segment anomalies; however, existing techniques are designed for industrial settings with a fixed camera position. In the built environment, images may be periodically collected by a moving camera manually or by aerial or ground vehicles and utilized for decision making. The camera pose between successive collections may vary widely voiding a fundamental assumption in existing anomaly detection approaches. To address this gap, we introduce the problem of Scene Anomaly Detection (Scene AD), where the goal is to detect anomalies from two sets of images: one set without anomalies and one set that may or may not contain anomalies. No labeled semantic segmentation data are provided for training. We propose a novel network, OmniAD, to tackle Scene AD by refining the reverse distillation anomaly detection method, leading to a 40% improvement in pixel-level anomaly detection. Additionally, we introduce two new data augmentation strategies that leverage novel view synthesis and camera localization to enhance generalization. We evaluate our approach both qualitatively and quantitatively on a new dataset, ToyCity the first Scene AD dataset featuring multiple objects as well as on the established single object centric dataset, MAD. Our method demonstrates marked improvement over baseline approaches, paving the way for robust anomaly detection in scenes with real-world camera pose variations commonly observed in the built environment. Project Page: https://drags99.github.io/OmniAD/
@@ 1 Introduction
Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (green), and query views which may contain anomalies (red). In contrast, traditional AD methods typically feature aligned views between the non-anomalous and anomalous images.
For many applications in the built environment —ranging from infrastructure assessment and construction monitoring to security surveillance and urban planning—detecting visual anomalies in scenes captured over time is a critical task  [1, 2, 3, 4, 5]. These anomalies may arise from structural changes, newly introduced objects, unexpected modifications in appearance, or more generally, anything that deviates from an established non-anomalous reference. However, images acquired at different instances, whether manually or autonomously, typically suffer from misalignment due to variations in camera pose. For example, bridge inspectors may assess changes to a structure following a disaster, self-driving car systems might detect modifications to traffic signs, and the U.S. Military places strong emphasis on monitoring infrastructure for anomalies critical to operational readiness [6, 7, 8, 9, 10, 11, 12]. In all such cases, each new set of images is unaligned with previous captures, complicating anomaly identification.
We define the task of Scene Anomaly Detection (Scene AD) as detecting anomalous regions in a query set of images, potentially containing no anomalies at all, against a non-anomalous reference set. Labeled examples of what qualifies as an anomaly are unavailable, rendering supervised segmentation methods impractical. Additionally anomalies are often too infrequent or unpredictable to compile sufficient labeled training data for every potential anomaly type, rendering fully supervised models infeasible. As a result, unsupervised anomaly detection (AD) approaches are a feasible alternative. However, these AD methods were originally crafted for controlled industrial scenarios which commonly assume fixed camera views and struggle with the viewpoint variations of real-world Scene AD.
Two main challenges emerge. First, there is no suitable multi-object, multi-view benchmark for Scene AD: most existing AD datasets capture a single view or focus on a single object [13]. Second, any robust Scene AD solution must tolerate significant camera misalignment and still detect newly introduced or changed regions.
To address these gaps, we propose OmniAD, a refinement of the Reverse Distillation (RD) framework [14]. RD detects anomalies by comparing a student decoder’s reconstruction with a teacher encoder’s embedding. We hypothesize that broadening the student decoder’s effective receptive field (ERF) [15] will improve sensitivity to anomalous regions. Accordingly, we incorporate student attention modules to expand the ERF and enhance pixel-level anomaly detection.
Moreover, to manage viewpoint differences in real scenes, we introduce two data augmentation strategies leveraging Neural Radiance Fields (NeRF) [16] and hierarchical localization (hloc) [17, 18]. These techniques generate novel, more closely aligned views of the query scene, aiding unsupervised methods—such as OmniAD—in detecting anomalies without labeled examples.
Since no comprehensive benchmark exists for this multi-object, multi-view setting, we introduce the ToyCity dataset. ToyCity consists of scenes with multiple objects, varied camera viewpoints, and different anomalies. We evaluate OmniAD and several baseline methods on ToyCity as well as on the single-object-centric Multi-pose Anomaly Detection dataset (MAD) [13], demonstrating notable performance gains.
@@
    Scene AD Task: Formal introduction of Scene Anomaly Detection, highlighting the challenge of detecting anomalies in unaligned multi-view imagery.
@@
    ToyCity Dataset: A new multi-object, multi-view dataset designed for benchmarking Scene AD methods in realistic settings.
@@
    OmniAD Architecture: A refined Reverse Distillation approach with a proposed Student Attention Module to expand the ERF, achieving state-of-the-art performance on ToyCity and MAD [13] and validated with ablation studies.
@@
    NeRF-based Data Augmentation Methods: Two novel techniques using NeRF and hierarchical localization for improved view-invariance, yielding a 40% increase in anomaly localization accuracy.
The rest of the paper is organized as follows: Section 2 reviews related work on anomaly detection, neural radiance fields, and camera localization. Section 3 introduces the proposed OmniAD architecture and novel data augmentation strategies. Section 4 presents the ToyCity dataset. Section 5 details the experimental setup, metrics, and results. Section 6 discusses the limitations of our dataset and method. Finally, Section 7 concludes the paper.
@@ 2 Related Work
Anomaly Detection. AD algorithms typically harness a collection of non-anomalous images , to establish a normative distribution. AD is a critical function in industrial visual inspection for identifying defects [19] and is also employed in healthcare for the detection of irregularities in medical images [20]. Prominent AD methods include CFLOW-AD [21], FastFlow [22], CFA [23], DRÆM [24], and Reverse Distillation (RD) [14].
@@
RD adopts a knowledge distillation strategy [26, 27], where a student decoder network learns the feature space representation from a teacher encoder network at various feature map scales. Unlike traditional approaches that feed the input image into both the student and the teacher networks, RD takes an autoencoder-like approach where the teacher encodes the input image, and the student decodes a fused multi-scale feature map output by the teacher. This feature map is a compressed representation of the teacher’s features across multiple scales.
Predominant datasets for benchmarking these methods, like MVTecAD [19] and VisA [28], typically employ a fixed overhead camera perspective, resulting in a uniform view of anomalies. Our research demonstrates that significantly different views can lead to substantial degradation in the pixel-level localization of anomalies. To combat these limitations, our method leverages a novel view synthesis method and camera localization techniques to generate corresponding non-anomalous images, thus facilitating improved anomaly detection. While there are existing methods for 3D anomaly detection, they generally require 3D data that is aligned with a reference model.
@@
Neural Radiance Field NeRFs, as introduced by Mildenhall et al. [16], represent a significant advancement in the field of view synthesis. These fields use a collection of images with known camera poses to synthesize new viewpoints of a scene. Unlike conventional 3D Structure-from-Motion (SfM) methods that construct explicit geometric models, NeRFs model a scene’s volume through a neural network that predicts color and density for space points along camera rays. This technique enables the creation of photorealistic images from novel perspectives. Subsequent innovations like Nerfacto [29] and Instant-ngp [30] have further refined NeRFs, offering rapid training and real-time rendering capabilities. In the context of anomaly detection, we leverage NeRFs to synthesize non-anomalous images from diverse viewpoints, significantly improving the detection process.
@@
Hierarchical Localization The process of hierarchical localization (hloc) [31, 18] has been developed to accurately determine a camera’s position and orientation in space, a task known as 6-Degree-of-Freedom (DoF) localization [31, 18, 32]. This process begins with the extraction of local features from a query image, utilizing methods such as SuperPoint [17], Sift [33], or DISK [34]. A 3D SfM model serves as a reference for these local features, against which the query image is matched using global descriptors like VLAD [35]. The Perspective-n-Point algorithm, in conjunction with RANSAC, then utilizes these correspondences to estimate the camera’s pose. This pose is critical in enabling NeRFs to generate non-anomalous images that correspond to those containing anomalies, thus in theory providing a more robust dataset for anomaly detection algorithms. However as we shall show an improved AD method is needed to fully utilize these anomaly aligned images.
@@ 3 Method
An overview of our AD methodology for Scene AD can be seen in Figure  2. Scene AD consists of non-anomalous views and query views. We generate an SfM model using non-anomalous views to acquire camera poses. Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure  3. For the first strategy, we interpolate between the non-anomalous camera poses. The second strategy for generating useful camera views for training is to utilize hloc to acquire the camera poses of query views. We then use NVS with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD. The following subsections describes each of these steps in more detail.
@@
Figure 3: Overview of our novel view selection strategy. Utilizing the non-anomalous dataset, we generate a structure from motion (SfM) 3D model and accompanying camera pose for each image. Leveraging these non-anomalous camera poses, we design a trajectory that connects them, allowing for interpolation. Subsequently, we apply a NVS technique to generate images that present unique perspectives along the interpolated path. The anomaly dataset can also be used to determine more novel views.
@@ 3.1 Novel View Selection Strategy
To generate new non-anomalous camera poses for NVS, we propose and evaluate two different approaches, namely Interpolated Non-anomalous Views (INV) and Query-aligned non-anomalous views (QANV).
3.1.1 Interpolated Non-anomalous Views
@@
Given a set of non-anomalous images an associated camera poses $T$ can be calculated using SfM. The calculated $T$ can be interconnected to form a trajectory. A simple approach for this interconnection is a greedy algorithm that selects a starting camera position and sequentially moves to the nearest neighboring camera position. Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles. An example of the camera interpolation for ToyCity can be seen in Figure  4.
3.1.2 Query-Aligned Non-anomalous Views
Here we propose a method to generate images of the baseline scene, but using poses from the anomaly scene. Our method employs hloc to accurately estimate the camera pose for a query anomaly image. This necessitates an accurate reconstruction of the baseline scene that is provided by the baseline SfM model. Hloc provides a variety of options for local feature detection, matching, and image similarity measurement. We opted for a combination of SuperPoint trained on Aachen and SuperGlue [17, 36], as it demonstrated effective performance on our datasets on average. Nonetheless, the flexibility to switch to other methods is inherent in our approach, enabling adaptation to specific dataset characteristics.

### v3

@@ Abstract
The built environment, encompassing critical infrastructure such as bridges and buildings, requires diligent monitoring of unexpected anomalies or deviations from a normal state in captured imagery. Anomaly detection methods could aid in automating this task; however, deploying anomaly detection effectively in such environments presents significant challenges that have not been evaluated before. These challenges include camera viewpoints that vary, the presence of multiple objects within a scene, and the absence of labeled anomaly data for training. To address these comprehensively, we introduce and formalize Scene Anomaly Detection (Scene AD) as the task of unsupervised, pixel-wise anomaly localization under these specific real-world conditions. Evaluating progress in Scene AD required the development of ToyCity, the first multi-object, multi-view real-image dataset, for unsupervised anomaly detection. Our initial evaluations using ToyCity revealed that established anomaly detection baselines struggle to achieve robust pixel-level localization. To address this, two data augmentation strategies were created to generate additional synthetic images of non-anomalous regions to enhance generalizability. However, the addition of these synthetic images alone only provided minor improvements. Thus, OmniAD, a refinement of the Reverse Distillation methodology, was created to establish a stronger baseline. Our experiments demonstrate that OmniAD, when used with augmented views, yields a 64.33% increase in pixel-wise $F_{1}$ score over Reverse Distillation with no augmentation. Collectively, this work offers the Scene AD task definition, the ToyCity benchmark, the view synthesis augmentation approaches, and the OmniAD method. Project Page: https://drags99.github.io/OmniAD/
@@ 1 Introduction
Detecting anomalies is critical for many scene assessment and analysis applications in the built environment, including infrastructure inspections[1, 2, 3], disaster response[4, 5], construction management [6] and planning [7]. Anomalies can be described as a deviation (e.g., damage, structural changes, new objects) from an established non-anomalous reference. In many scene assessment and analysis applications, images are the primary sensing modality, as they are inexpensive to capture and provide the dense resolution required for visual evidence in documentation [8, 9].
Anomaly detection (AD) approaches could be categorized as supervised or unsupervised. Supervised approaches rely on annotated data and work well when high-quality annotations are available but can be impractical in many real-world inspection scenarios where anomalies are rare or unknown in advance [10, 11, 12]. Unsupervised AD methods, on the other hand, are often more practical as they don’t require labeled data of anomalies. Current benchmarks for unsupervised anomaly detection have largely been tailored for manufacturing contexts which consist of imagery of a single-object [13, 14, 12] or multiple objects [15, 16] with constrained views.
While unsupervised image-based AD methods have found great success in manufacturing applications, they have not been explored for scene analysis applications in the built environment, where many manufacturing-specific assumptions no longer hold. Imagery in the built environment is typically collected from unconstrained viewpoints, often by handheld or UAV-mounted cameras [1, 17, 18, 19]. As shown in Figure 1 view-points for images captured in the built environment vary in nature tremendously, unlike the fixed-camera setting typically used for unsupervised AD benchmarks [13, 12, 20, 15]. Seminal anomaly detection approaches [21, 22, 23, 24, 25] discussed in more detail in the related works section have matured to the point of achieving very high quantitative metrics with fixed views, with state-of-the-art methods plateauing in performance. However, our studies showed that seminal works are unable to meet similar performance levels for multi-view multi-object scenes such as in Figure 1 (a). This performance disparity underscores the need to specifically define and address the problem of anomaly detection within these unconstrained visual environments.
An effective scene anomaly detection method in unconstrained visual environments must operate (i) without labeled data, (ii) across multiple objects per image, and (iii) under random viewpoint variation. We formally refer to this general AD task as Scene Anomaly Detection (Scene AD) and illustrate it in Figure 2. Formally, Scene AD entails localizing anomalous regions given two unlabeled image sets: one depicting normal reference conditions, called the non-anomalous set, and one containing potential anomalies, if any, called the query set. No assumption is made for alignment between the non-anomalous and query set. Camera poses, if needed, should only be estimates from methods such as structure-from-motion or GPS measurements to mimic the noisy camera poses that can feasibly be obtained in an uncontrolled environment.
To provide a benchmark for Scene AD, we introduce ToyCity a real-world dataset consisting of four multi-object scenes, with more than 1200 anomalous images per scene and over 20 varying anomalies across scenes. The diversity of anomalies and quantity of anomalous images in ToyCity match or exceed those of the most frequently used fixed view AD benchmarks [15, 26, 16, 27, 13, 20, 14]. ToyCity also exceeds the real-image diversity of the Multi-Pose Anomaly Detection (MAD)[12] dataset, which contains 16 to 32 anomalous images per object. When evaluating baseline AD methods, that perform well on prior datasets, we found they exhibit very weak pixel-level localization on ToyCity. Among the baseline AD methods, Reverse Distillation (RD) [25] showed the best localization performance, however we found its localization performance to be inconsistent for practical use.
We introduce OmniAD to validate the potential of unsupervised AD approaches for pixel-level Scene Anomaly Detection. OmniAD is a refinement of Reverse Distillation [25] that integrates student-attention modules. These modules are designed to expand the student decoder’s effective receptive field (ERF) [28], thereby improving its sensitivity to localized anomalies. Furthermore, to counteract the performance degradation caused by viewpoint variations in Scene AD, we explore two data augmentation strategies. Both strategies aim to generate novel non-anomalous views, using Neural Radiance Fields (NeRF) [29] and hierarchical localization (hloc) [30, 31], to assist AD methods in generalizing to varying views.
Figure 1: (a) Typical images used in the built environment. (b) Images available for benchmarking unsupervised anomaly detection [13, 15, 16]. The top row for (a) and (b) consist of non-anomalous images, while the bottom is anomalous.
@@
Figure 2: In Scene AD, we observe variations in camera positions between non-anomalous views (green) and query views that may contain anomalies (red). In contrast, traditional AD methods typically feature aligned views between the non-anomalous and anomalous images.
@@
    Scene AD Task: Formal introduction of Scene Anomaly Detection, highlighting the challenge of detecting anomalies in unaligned multi-view imagery consisting of multiple objects.
@@
    ToyCity Dataset: A new multi-object multi-view dataset of real images designed for benchmarking methods for Scene AD.
@@
    Baselines: We establish the first quantitative metrics of seminal AD methods for Scene AD using ToyCity.
@@
    NeRF based Data Augmentation Methods: Two augmentation strategies for increased view-generalization in Scene AD.
  • 5.
    OmniAD: A refinement of Reverse Distillation [25], which when combined with augmentation strategies leads to a 64.33% improvement in pixel-wise $F_{1}$ score for Scene AD compared to Reverse Distillation with no augmentation.
The rest of the paper is organized as follows: Section 2 reviews related work on anomaly detection, neural radiance fields, and camera localization. Section 3 introduces the proposed OmniAD architecture and novel data augmentation strategies. Section 4 presents the ToyCity dataset. Section 5 details the experimental setup, metrics, and results. Section 6 discusses the limitations of our dataset and method. Finally, Section 7 summarizes and concludes this paper.
@@ 2 Related Work
2.1 Unsupervised Image-Based Anomaly Detection
Anomaly detection (AD) algorithms typically harness a collection of non-anomalous images, to establish a normative distribution. AD is a critical function in industrial visual inspection for identifying defects [10] and is also employed in healthcare for the detection of irregularities in medical images [32]. Unsupervised AD methods have demonstrated considerable success, particularly for fixed-view single-object anomaly detection. This domain has become mature over time to achieve a high-performance plateau for pixel-wise anomaly detection. New methods [33, 34, 35, 36] frequently benchmark against a core group of foundational techniques [21, 22, 23, 24, 25] due to the high-performance plateau achieved.
When introducing novel tasks such as Scene AD, we must select robust baseline methods that are demonstrably reproducible and widely accepted for comparison to create a meaningful benchmark. Accordingly, our analysis centers on prominent methods such as CFLOW-AD [21], FastFlow [22], Coupled-hypersphere-based Feature Adaptation (CFA) [23], discriminatively trained reconstruction anomaly embedding model (DRÆM) [24], and Reverse Distillation (RD) [25]. These specific methods cover the two primary categories unsupervised image-based AD methods fall into, feature embedding based and reconstruction based [12, 34]. Furthermore, this strategic selection aims to establish a durable and widely understood performance benchmark. These methods offer strong guarantees of reproducibility, stemming from their continuous validation and the availability of robust, often multiple, implementations [37].
2.2 Feature Embedding Based Unsupervised Anomaly Detection
@@ 2.3 Reconstruction Based Unsupervised Anomaly Detection
Reverse Distillation adopts a knowledge distillation strategy [39, 40], where a student decoder network learns the feature space representation from a teacher encoder network at various feature map scales. Unlike traditional approaches that feed the input image into both the student and the teacher networks, Reverse Distillation takes an autoencoder-like approach where the teacher encodes the input image, and the student decodes a fused multiscale feature map output by the teacher. Deviations between the student and teachers’ features indicate the occurrence of an anomaly.
2.4 Multi-View Anomaly Detection
While the aforementioned unsupervised AD methods demonstrate strong performance in controlled environments, a significant limitation inherent in many of these 2D approaches arises when deployed in multi-view settings [12]. Their effectiveness often depends critically on feature representations or learned normality models (such as distributions, reconstruction targets, or feature manifolds) that are derived from training data [10, 11]. As a result, these methods frequently struggle when presented with the significant viewpoint variations from those seen in the training data that is typical of real-world data acquisition, like images captured using handheld devices or UAVs during infrastructure assessments [1, 19]. Additionally, realistic scenes often contain multiple objects simultaneously, a scenario largely distinct from typical single-object AD benchmarks. Consequently, the specific task of unsupervised anomaly detection within environments characterized by both multiple objects per scene and random, unconstrained multi-view captures has remained largely unexplored. This is partly due to the lack of suitable benchmark datasets designed to address these combined complexities, as the manual effort in creating such a dataset with real images is tremendous in comparison to labeling the occurrence of an anomaly on a single object.
2.5 Neural Radiance Field
@@
Neural Radiance Fields (NeRFs) [29] represent a significant advancement in the field of view synthesis. These fields use a collection of images with known camera poses to synthesize new viewpoints of a scene. Unlike conventional 3D Structure-from-Motion (SfM) methods that construct explicit geometric models, NeRFs model a scene’s volume through a neural network that predicts color and density for space points along camera rays. This technique enables the creation of photorealistic images from novel perspectives. Subsequent innovations like Nerfacto [41] and Instant-ngp [42] have further refined NeRFs, offering rapid training and real-time rendering capabilities. In the context of this work, we leverage NeRFs to synthesize non-anomalous images from diverse viewpoints.
2.6 Hierarchical Localization
@@
The process of hierarchical localization (hloc) [43, 31] has been developed to accurately determine a camera’s position and orientation in space, a task known as 6-Degree-of-Freedom localization [43, 31, 44]. This process begins with the extraction of local features from a query image, utilizing methods such as SuperPoint [30], Sift [45], or DISK [46]. A 3D SfM model serves as a reference for these local features, against which the query image is matched using global descriptors like VLAD [47]. The Perspective-n-Point algorithm, with RANSAC, then utilizes these correspondences to estimate the camera’s pose. This pose is critical in enabling NeRFs to generate non-anomalous images that correspond to those containing anomalies, thus in theory providing a more robust dataset for anomaly detection algorithms. However, as we shall show, an improved AD method is needed to fully utilize these anomaly aligned images.
@@ 3 Method
An overview of our AD methodology for Scene AD can be seen in Figure 3. Scene AD consists of non-anomalous views and query views. We generate an SfM model using non-anomalous views to acquire camera poses. Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation, as shown in Figure 4. For the first strategy, we interpolate between the non-anomalous camera poses. The second strategy for generating useful camera views for training is to utilize hloc to acquire the camera poses of query views. We then use a NeRF for novel view synthesis (NVS) with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD. The following subsections describe each of these steps in more detail.
@@
Figure 4: Overview of our novel view selection strategy. Utilizing the non-anomalous dataset, we generate a structure from motion (SfM) 3D model and accompanying camera pose for each image. Leveraging these non-anomalous camera poses, we design a trajectory that connects them, allowing for interpolation. Subsequently, we utilize a NeRF for Novel View Synthesis (NVS) to generate images that present unique perspectives along the interpolated path. Query images can also be used to determine more novel views.
@@ 3.1 Novel View Selection Strategy
A core challenge in Scene AD is the viewpoint variation between the non-anomalous reference images and the query images. Anomaly detection models trained only on the original non-anomalous views may struggle when presented with a query image from a significantly different perspective. To enhance the model’s robustness to these variations, a wider variety of non-anomalous viewpoints are required for training, ideally including views similar to those encountered in the query set.
To generate these new non-anomalous camera poses for novel view synthesis (NVS), we propose and evaluate two different approaches:
  • 1.
    Interpolated Non-anomalous Views (INV): The goal of this strategy is to densify the coverage of the non-anomalous viewpoint space. By generating views between the existing captured non-anomalous poses, we train the model to handle gradual shifts in perspective, improving its general robustness to unseen viewpoints near the original capture trajectory.
  • 2.
    Query-Aligned Non-anomalous Views (QANV): This strategy directly aims to address the misalignment with potentially anomalous query images. By estimating the camera pose of a query image and then synthesizing a non-anomalous view from that approximate camera pose, we provide the model with similar normal/potentially anomalous pairs (in terms of viewpoint). Intuitively, this allows the model to better isolate deviations caused by anomalies rather than viewpoint changes.
The following subsections detail how these camera poses are generated and used with NVS.
3.1.1 Interpolated Non-anomalous Views (INV)
Given a set of non-anomalous images, an associated camera poses $T$ can be calculated using SfM. The calculated $T$ can be interconnected to form a trajectory. A simple approach for this interconnection is a greedy algorithm that selects a starting camera position and sequentially moves to the nearest neighboring camera position. Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles. An example of the camera interpolation for ToyCity can be seen in Figure  5.
@@ 3.1.1 Interpolated Non-anomalous Views (INV)
3.1.2 Query-Aligned Non-anomalous Views (QANV)
Here, we propose a method to generate additional camera poses relative to the non-anomalous scene by localizing images from the query set. Our method employs hloc to estimate the camera pose for a query anomaly image. This necessitates an accurate reconstruction of the non-anomalous scene that is provided by the non-anomalous SfM model. Hloc provides various options for local feature detection, matching, and image similarity measurement. We opted for a combination of SuperPoint trained on Aachen and SuperGlue [30, 48], as it demonstrated effective performance on our datasets on average. Nonetheless, the flexibility to switch to other methods is inherent in our approach, enabling adaptation to specific dataset characteristics.